Mathematics 242 – Principles of Analysis Definitions Quiz 7 – April 12, 2013

Your Name:

A) (5) What does it mean to say that a set is countably infinite?

B) (5) Let

f(x) = \{ \begin{matrix}1 − 3x & if x < 0 \\ 1 + 2x & if x ≥ 0. \end{matrix}

Is f differentiable at c = 0? Why or why not?